OpenStudy (ebydaqueen19):

help!!! 4radical81a8b8 simplify

OpenStudy (anonymous):

May i know what is a and b here?? 4radical 81 is 3

OpenStudy (ebydaqueen19):

its a to the power of 8 and b to the power of 8

OpenStudy (ebydaqueen19):

is tht what u mean?

OpenStudy (zzr0ck3r):

\(\sqrt[4]{81a^8b^8}\) \(\large ?\)

OpenStudy (zzr0ck3r):

\(\sqrt[4]{81a^8b^8}=\sqrt[4]{81}\sqrt[4]{a^8}\sqrt[4]{b^8}=3a^{\frac{8}{4}}b^{\frac{8}{4}}=3a^2b^2\)

OpenStudy (ebydaqueen19):

thank you so much

OpenStudy (zzr0ck3r):

notice that sometimes its nicer to think about the radical form and sometimes its nice to work with the fractional exponent form

OpenStudy (zzr0ck3r):

but over and over and over you will use \(\sqrt[c]{ab}=\sqrt[c]{a}\sqrt[c]{b}\)